# Training Software Engineering Agents and Verifiers with SWE-Gym

### v1

@@ 3.2 Training General-Purpose Prompting Agents
[…]
@@
7B $40.3$ $29.7$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}-10.7}$ $47.0$ $31.0$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}-16.0}$ $20.3$ $22.2$ $+1.9$ $1.0(\pm 1.0)$ $10.0(\pm 2.4)$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+9.0}$
14B $49.7$ $\mathbf{18.1}$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}-31.6}$ $31.7$ $27.1$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}-4.6}$ $23.2$ $21.4$ $-1.8$ $2.7(\pm 1.9)$ $12.7(\pm 2.3)$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+10.0}$
32B $\mathbf{27.0}$ $\mathbf{18.1}$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}-8.9}$ $\mathbf{16.7}$ $\mathbf{18.1}$ ${\color[rgb]{0.86328125,0.078125,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.078125,0.234375}+1.5}$ $\mathbf{15.5}$ $29.3$ $+13.9$ $\mathbf{3.0}(\pm 1.4)$ $\mathbf{15.3}(\pm 2.5)$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\mathbf{+12.3}}$
@@
7B $45.8$ $33.8$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}-12.0}$ $39.6$ $\mathbf{21.0}$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}-18.6}$ $21.9$ $35.3$ $+13.4$ $1.8(\pm 1.1)$ $10.6(\pm 2.1)$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+8.8}$
14B $44.9$ $14.5$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}-30.4}$ $32.1$ $21.3$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}-10.7}$ $25.5$ $30.1$ $+4.6$ $4.0(\pm 1.6)$ $16.4(\pm 2.0)$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+12.4}$
32B $\mathbf{9.5}$ $\mathbf{13.8}$ ${\color[rgb]{0.86328125,0.078125,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.078125,0.234375}+4.3}$ $\mathbf{29.4}$ $23.8$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}-5.6}$ $24.6$ $31.6$ $+7.0$ $\mathbf{7.0}(\pm 1.3)$ $\mathbf{20.6}(\pm 2.1)$ ${\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\mathbf{+13.6}}$
Table 3: Model performance (fine-tuned on 491 SWE-Gym-sampled trajectories) on SWE-Bench (Jimenez et al., 2024) using OpenHands (Wang et al., 2024c) as agent scaffold. We use Qwen-2.5-Coder-Instruct as the base model. We set temperature $t=0$ for evaluation.
@@
In this subsection, we use OpenHands (version CodeActAgent 2.1, Wang et al. 2024c, b) as our agent scaffold. OpenHands CodeActAgent uses a ReAct-style (Yao et al., 2023) general-purpose prompting without specialized workflow5.1) and relies on the underlying LM to perform action and interpret observation. It equips an LM with a bash terminal and a file editor. We disable the browser feature of OpenHands in this work. We use OpenHands’s remote runtime (Neubig & Wang, 2024) feature to roll out agents in parallel (i.e., execute agent’s actions) on SWE-Gym.
@@
Trajectory Collection. We roll out 491 successful trajectories where the agent successfully solves the given task from SWE-Gym. These trajectories are sampled from two models (gpt-4o-2024-08-06 and claude-3-5-sonnet-20241022) with different temperature settings. Each success trajectory, on average, has 39.9 LM completion messages (roughly 19 turns) and $18,578$ tokens. Please refer to Tab. 8 for more details. While SWE-Gym contains many more tasks for trajectory collection and allows repeated sampling, our current set of 491 trajectories was primarily limited by computational budget constraints rather than the availability of tasks. Please refer to Tab. 10 and §B.3 for details on the temperature setting and the resolve rate in the training set.
Additional Evaluation Metrics & Settings. In addition to the metrics presented in §3.1, we include the following for further analysis: (3) Stuck in Loop (%), the percentage of trajectories where the agent is stuck in a loop by repeating the exact same action for the last three turns; (4) Avg. Turn(s), the average number of action-observation turns for these trajectories. Our evaluation of the trained LM is bounded by either 100 interaction turns or the 32k context window length, whichever is reached first. We use sampling temperature of 0 (i.e., greedy) unless otherwise specified.
[…]
@@
Training help reduce agent’s stuck-in-loop behavior. Open-weight LMs, often suffer from stuck-in-loop, especially when prompted with general-purpose prompts (§5.1): the LM would repeat exactly the same action for multiple turns and never get out of this loop. This is evident in the zero-shot stuck in loop statistics in Tab. 3, where even the largest 32B model gets stuck in loop $30\%$ of the time. The fine-tuned model consistently reduces the stuck in loop rate ranging from $-4.6\%$ to $-18.6\%$ for problems in both SWE-Bench Lite and Verified, with the exception of 32B model on SWE-Bench Lite ($+1.5\%$). The reduced stuck-in-loop rate likely encourages the agent to execute more code edits, as evident by the decreased empty patch rate across different scales (except for 32B on SWE-Bench Verified).
@@
Performance scales with model size. Rather unsurprisingly, we see the resolve rate, the empty patch, and the stuck-in-loop rate improves consistently in both SWE-Bench Lite and Verified as we switch to bigger base models (Tab. 3).
Self-improvement is not yet working. We use the fine-tuned 32B model to rollout trajectories from SWE-Gym for 6 times and obtain 868 successful trajectories (i.e. on-policy trajectories) using temperature $t=0.5$. We further fine-tune the base 32B model on a mixture of 868 on-policy trajectories and the previously collected 491 off-policy trajectories. When evaluating the model on SWE-Bench Lite, we observe a performance drop from 15.3% to 8.7% compared to the model that was fine-tuned using only off-policy trajectories from GPT-4 and Claude.
We hypothesize improved results could be achieved with more advanced optimization methods, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), or by employing a stronger base model. These directions remain promising avenues for future investigation.
3.3 Self-improvement with Specialized Workflow Agents
@@
We then explore the opposite end of the scaffold spectrum—agents with specialized workflows—using the MoatlessTools Agent (version 0.0.2; Örwall 2024) for our experiments. Unlike OpenHands CodeAct, which offers extensive freedom in long-horizon planning, this scaffold constrains the model’s action space with pre-defined workflows, effectively reducing task horizons. As demonstrated in Tab. 3 and Tab. 5, we observe that open-weight models consistently deliver better zero-shot performance when paired with the MoatlessTools scaffold.
@@
Given the improved zero-shot performance and shorter task horizon with the MoatlessTools scaffold, we explore if SWE-Gym allows agents to self-improve without reliance on a strong teacher. With a limited compute budget, we conduct this experiment with only 7B and 34B models, using LoRA (Hu et al., 2022) for the 34B models for improved efficiency.
We start our experiments with ablations on 7B model and scale it to 32B. We sample the model for 30 rounds on SWE-Gym-Lite with a high temperature of 1.0, adding successful trajectories to the fine-tuning dataset. This process is repeated twice, after which the improvements are marginal.
@@
Easy Data Bias Degrades Model Performance. During repeated sampling, similar to the observation in  Brown et al. (2024) , we find that the success probability for each instance follows a long-tail distribution as shown in Fig. 3, with more instances being solved simply by increasing the number of samples. While having a broader coverage of tasks should be beneficial for training, as first observed in math reasoning (Tong et al., 2024), repeated sampling introduces a distribution bias toward easier tasks, making it suboptimal to naively train on all successful trajectories.
Mitigate Easy Data Bias with Per Instance Capping. To mitigate such bias, Tong et al. (2024) propose to keep the training set size same but obtain equal or more number of positive instances on difficult tasks. However, obtaining positive instances from harder tasks is inherently difficult and on average requires more rounds of sampling and therefore more compute. We consider a compute-bounded scenario with a fixed sampling budget to optimize model performance through data selection.
We propose per-instance capping - a simple method limiting the maximum instances per task. As shown in Fig. 3, this approach balances dataset bias and size. Too low a cap limits dataset size and hurts performance (see §4.2), while too high a cap skews distribution toward easier tasks.
As shown in Tab. 4, empirically, we find a threshold of 2 achieves a good balance between dataset size and difficulty bias, performing marginally better than the full dataset while improving training speed. We rank trajectories by the rounds of model responses they need, fewer the better.
[…]
0 (Zero-shot) 0 56.3 7.0
1 36 37.3 9.0
2 62 29 9.7
3 82 43.7 7.7
No Cap (All) 172 30.7 9.3
Table 4: Resolve rate and empty patch rate on SWE-Bench Lite after 7B model trained with with data from different instance capping strategies (Cap) and therefore different number of trajectories (Traj).
@@
Table 5: Resolve rate (RR) and Empty patch rate (EP) on SWE-Bench Lite with MoatlessTools Scaffold after online rejection sampling fine-tuning, evaluated at temperature $t=0$. RR shown in highlighted columns.
Results. With a capping of 2, we perform two rounds of policy improvements for both 7B and 32B models and present results in Tab. 5. For the 7B model, the resolve rate improves significantly, rising from 7.0% in the zero-shot setting to 9.0% after the first iteration and reaching 10.0% after the second iteration, indicating notable benefits from training on SWE-Gym. In contrast, the 32B model demonstrates strong initial performance with a zero-shot resolve rate of 19.0%, but shows only marginal improvement to 19.7% after the first iteration and no additional gains in subsequent iterations. We hypothesize that this plateau may stem from limitations inherent to the agent scaffold and the rejection sampling finetuning algorithm we employed to improve the policy.
4 Scaling Agent Improvements with SWE-Gym
4.1 Inference-Time Scaling with SWE-Gym-powered Verifier
@@
Beyond training the agent, trajectories sampled from SWE-Gym also allow us to train a verifier (a.k.a. reward model). As a baseline, we train an outcome-supervised reward model (ORM) (Cobbe et al., 2021) that, given the relevant context of the task execution (e.g., problem statement, trajectories, git diff), generates a score representing the probability if the agent solves the problem.
We show that such learned verifiers enable effective inference-time scaling for further performance improvement. Using the same underlying LM, one can sample multiple solutions for the same problem and use the verifier to pick the best solution.
4.1.1 Verifier for Agent Scaffold with General-Purpose Prompting
For OpenHands CodeActAgent (Wang et al., 2024c, b) that uses general-purpose prompting (§5.1), we consider a verifier (ORM) setting where the input is an interleaved trajectory $\tau$ and output is a scalar reward $r$:
$\tau=[o_{1},a_{1},o_{2},a_{2},\dots,o_{n},a_{n}],r\in[0,1]$
where observation $o_{k}$ can be the problem-statement, command execution output, error messages, etc; action $a_{k}$ can be bash command or file operations (e.g., edit, view) from the agent. We include verifier prompt template in §B.5.
Training Setup. Following the same training hyper-parameters as described in §3, we train a 32B Qwen2.5-Coder-Instruct as verifier. We train the LM to predict ‘YES’ or ‘NO’ regarding if the agent has successfully solved the request based on the agent trajectory (see the verifier prompt in §B.5). At inference time, conditioned on the prompt and the agent trajectory $\tau$, we use SGLang (Zheng et al., 2024a) to obtain the log probability of the next token being ‘YES’ ($l_{y}$) or ‘NO’ ($l_{n}$). We then calculate the probability of success as ORM by normalizing the log probability: $p_{\text{yes}}=\exp(l_{y})/(\exp(l_{y})+\exp(l_{n}))$.
@@ 4.1.1 Verifier for Agent Scaffold with General-Purpose Prompting
Verifier Training Data. We re-use two sets of trajectories we sampled for agent training in §3.2: (1) off-policy trajectories which contain 443 successful trajectories that are sampled from gpt-4o and claude-3-5-sonnet; (2) on-policy trajectories which contain 875 successful trajectories sampled from the agent model (fine-tuned Qwen2.5-Coder-Instruct-32B). Note that we only keep trajectories that fit in the context window (32k tokens) for training. We combine both on-policy and off-policy trajectories, randomly sample the same amount of unsuccessful trajectories from each subset (1318 in total), and combine them as our dataset for verifier training.
Metrics. We measure (1) Pass@K, which represents the percentage of tasks where at least one successful solution is found among K sampled trajectories, and (2) Best@K, which selects the trajectory with the highest verification score ($p_{\text{yes}}$) among K samples and reports the percentage of these selected trajectories that are successful. Pass@K evaluates the model’s ability to find any working solution (i.e., the upper bound for Best@K), while Best@K assesses our verifier’s capability to identify the most promising solution. We calculate the mean and variance for each data point following Lightman et al. (2023), please see §B.1 for details on the calculation.
@@
Figure 4: Scaling inference-time compute improves performance on SWE-Bench Verified using a fine-tuned verifier. Both the agent and the verifier are Qwen2.5-Coder-Instruct-32B model fine-tuned on the corresponding dataset (§4.1.1). OpenHands (Wang et al., 2024c) is used as the agent scaffold. The first rollout was performed with temperature $t=0$, and $t=0.5$ was used for the rest.
@@
Result. Fig. 4 shows how Pass@K and Best@K metrics scale with the number of sampled agent trajectories using the fine-tuned 32B model as the agent model. Pass@K demonstrates strong improvement, rising from $20.6\%$ to $37.8\%$ as K increases from 1 to 8, and to $42.8\%$ at k=16. The Best@K metric, which relies on our verifier’s ability to select the best trajectory, demonstrates more modest but steady progress, improving from $20.6\%$ at k=1 to $29.8\%$ at k=8, and to $32.0\%$ at k=16. The gap between Pass@K and Best@K reveals room for more advanced reward modeling techniques for coding agents. Surprisingly, we found that fine-tuning the model using LoRA (Hu et al., 2022) (29.8% @$8$) via Unsloth library (Unsloth Team, 2024) performs better than full-parameter fine-tuning for verifier training (27.2%@$8$), potentially due to the effect of regularization. Furthermore, as shown in Fig. 1 (bottom), the Best@K curve exhibits strong linearity on a logarithmic scale, indicating a promising scaling behavior.
Figure 5: Abaltion study for verifier training (§4.1.1). Performances are evaluated on SWE-Bench Verified. Both the agent and the verifier are Qwen2.5-Coder-Instruct-32B model fine-tuned on the corresponding dataset (§4.1.1). OpenHands (Wang et al., 2024c) is used as the agent scaffold.
@@
Ablation Study: Training Data matters for Verifier. As shown in Fig. 5, our ablation study demonstrates that the choice of training data can significantly impact verifier performance. Different from the LoRA-trained model in the previous paragraph, all ablation studies perform full-parameter fine-tuning on a 32B Qwen2.5-Coder-Instruct for the verifier. Training with a mixture of off-policy and on-policy data yields the best results (our default setting), with Best@k scaling from $20\%$ to approximately $27\%$ at k=8. In contrast, using only on-policy data from the fine-tuned model shows moderate but limited improvement, while training exclusively on off-policy data from Claude and GPT leads to early performance plateaus around $22\%$. Notably, using off-policy data with twice the number of negative examples performs slightly worse with a lower number of $k$, suggesting that overweighting negative examples during training could make the verifier overly conservative. These findings indicate that verifier training benefits most from a diverse dataset combining both off-policy and on-policy examples, as well as balanced positive and negative data, enabling better generalization and more accurate identification of successful solutions during inference time scaling.
4.1.2 Verifier for Agent Scaffold with Specialized Workflow

### v2

@@ 4.2 Training General-Purpose Prompting Agents
Table 3: Model performance (fine-tuned on 491 SWE-Gym-sampled trajectories) on SWE-Bench (Jimenez et al., 2024) using OpenHands (Wang et al., 2024c) as agent scaffold. We use Qwen-2.5-Coder-Instruct as the base model.
Model Empty Patch (%, $\downarrow$) Stuck in Loop (%, $\downarrow$) Avg. Turn(s) Resolve Rate (%, $\uparrow$)
@@
7B 40.3 29.7 -10.7 47.0 31.0 -16.0 20.3 22.2 +1.9 1.0 ($\pm$ 1.0) 10.0 ($\pm$ 2.4) +9.0
14B 49.7 18.1 -31.6 31.7 27.1 -4.6 23.2 21.4 -1.8 2.7 ($\pm$ 1.9) 12.7 ($\pm$ 2.3) +10.0
32B 27.0 18.1 -8.9 16.7 18.1 +1.5 15.5 29.3 +13.9 3.0 ($\pm$ 1.4) 15.3 ($\pm$ 2.5) +12.3
@@
7B 45.8 33.8 -12.0 39.6 21.0 -18.6 21.9 35.3 +13.4 1.8 ($\pm$ 1.1) 10.6 ($\pm$ 2.1) +8.8
14B 44.9 14.5 -30.4 32.1 21.3 -10.7 25.5 30.1 +4.6 4.0 ($\pm$ 1.6) 16.4 ($\pm$ 2.0) +12.4
32B 9.5 13.8 +4.3 29.4 23.8 -5.6 24.6 31.6 +7.0 7.0 ($\pm$ 1.3) 20.6 ($\pm$ 2.1) +13.6
@@
In this section, we use OpenHands (version CodeActAgent 2.1, Wang et al. 2024b, c) as our agent scaffold, which is based on general-purpose ReAct-style prompting (Yao et al., 2023). In contrast to specialized-workflows-agents (§2), it relies on the LM to generate actions and do planning. It equips the base LM with a bash terminal and a file editor. We disable the browser feature of OpenHands in this work.
@@
Trajectory Collection. By rejection sampling, we obtain 491 successful trajectories from SWE-Gym,. These trajectories are sampled from gpt-4o-2024-08-06 and claude-3-5-sonnet-20241022 with different temperature settings. Each successful trajectory, on average, has roughly 19 turns and approximately 19,000tokens. [^5] Tab. 8 contains more statistics of the sampled trajectories. Although SWE-Gym offers many more tasks and allows repeated sampling, our 491 trajectories are limited primarily by computational budget.
Training on SWE-Gym trajectories turns LM into effective agents to fix issues. As shown in Tab. 3, the pre-trained base model achieves resolution rates of 3.0% and 7.0% on SWE-Bench Lite and Verified, respectively. After fine-tuning on 491 trajectories [^6] We use a sampling temperature of 0 unless otherwise specified., it improves by up to 12.3% (3.0% → 15.3%) and 13.6% (7.0% → 20.6%).
@@
Training reduces stuck-in-loop behavior. For agent tasks, open-weight LMs often get stuck in loops, where the model perpetually generates the same action for multiple turns, especially when prompted with general-purpose prompts (§2). Thus, we report Stuck in Loop (%), the percentage of trajectories where the agent repeats the same action three times consecutively. As shown in Tab. 3, zero-shot pre-trained models often get stuck in loops; even the largest 32B model is trapped in 29.4% of SWE-Bench Verified tasks. Fine-tuning on trajectories from SWE-Gym consistently reduces the stuck-in-loop rate by 4.6–18.6% across both SWE-Bench Lite and Verified tasks, except for the 32B model on SWE-Bench Lite, which increases by 1.5% due to its already low loop rate. This coincides with a decrease in the empty patch rate, likely enabling the agent to perform more code edits.
@@
Performance scales with model size. Rather unsurprisingly, larger base models consistently improve the resolution rate, empty patch rate, and stuck-in-loop rate (Tab. 3).
Self-improvement remains ineffective. In addition to fine-tuning on trajectories sampled from strong teacher models, we also experiment with fine-tuning on trajectories sampled directly from the policy being updated. We use the fine-tuned 32B model to sample 6 trajectories per SWE-Gym instance (using temperature $t=0.5$), obtaining 868 successful trajectories (i.e., on-policy trajectories). We further fine-tune the base 32B model on a mixture of 868 on-policy trajectories and the previously collected 491 off-policy trajectories. When evaluating this fine-tuned model on SWE-Bench Lite, we observe the resolution rate drop from 15.3 to 8.7%, suggesting that self-improvement is not yet working. We hypothesize that we could achieve improved results using more advanced policy optimization methods, such as proximal policy optimization (PPO) (Schulman et al., 2017), or with a stronger base model. These directions remain promising avenues for future investigation.
4.3 Self-Improvement with Specialized Workflow
@@
Unlike OpenHands, which offers freedom in long-horizon planning, MoatlessTools constrains the language model’s action space to pre-defined specialized workflows, reducing task horizons. Specialized workflows outperform general-purpose prompting for open-weight LMs. In Tab. 3 and Tab. 4, the 7B and 32B LM achieve zero-shot resolution rates of 7% and 19% with MoatlessTools, compared to 1.0% and 3.0% with OpenHands on SWE-Bench Lite.
@@
Given MoatlessTools’ improved zero-shot performance and shorter task horizon, we hypothesize that self-improvement without a strong teacher is achievable using this scaffold and training on SWE-Gym. With a limited compute budget, we conduct this experiment with only 7B and 32B models, using LoRA (Hu et al., 2022) for the 32B model for improved efficiency. We use the 7B model for ablation experiments.
We use iterative rejection sampling fine-tuning for policy improvement. Each iteration involves (a) performing 30 high-temperature (1.0) rollouts per task on SWE-Gym-Lite and adding successful trajectories to the fine-tuning dataset, and (b) fine-tuning the policy on these filtered trajectories. After two iterations, further improvements are negligible.
@@
Data Bias Impacts Performance. Repeated sampling, as in Brown et al. (2024), shows that task success probability follows a long-tail distribution (Fig. 6), where more samples increase solved instances. While broader task coverage benefits training, it introduces a bias toward easier tasks, making it suboptimal to train on all successful trajectories, as first observed in math reasoning Tong et al. (2024).
Mitigating Bias with Per-Instance Capping. We introduce per-instance capping—a method that limits the maximum number of selected samples per task. As illustrated in Fig. 6, this balances dataset bias and size. A low cap reduces dataset size and performance (§5.2), while a high cap skews the distribution toward easier tasks. Empirically, a threshold of 2 achieves a good balance, slightly outperforming the full dataset and improving training speed (Tab. 6). We rank trajectories by the number of model response rounds required, preferring fewer.
Results. Results. After two policy improvement iterations (Tab. 4), the 7B model’s resolution rate increased from 7.0% to 9.0% after the first iteration and to 10.0% after the second. In contrast, the 32B model improved from 19.0% to 19.7% after the first iteration with no further gains. We attribute the limited gains in the 32B model to the scaffold’s restricted action space and the rejection sampling fine-tuning method.
Table 4: resolution rate (RR) and Empty patch rate (EP) on SWE-Bench Lite with the MoatlessTools Scaffold after online rejection sampling fine-tuning (temperature $t=0$).
@@
5 Scaling Agent Performance with SWE-Gym
We explore two scaling directions enabled by SWE-Gym to enhance agent performance: inference-time scaling (§5.1) and training-time data scaling (§5.2).
5.1 Inference-Time Scaling with Verifiers
@@
Trajectories sampled from SWE-Gym can be used not only for training a policy, but also for training a verifier (i.e., reward) model. We train an outcome-supervised reward model (ORM) (Cobbe et al., 2021) that, given the relevant context of the task execution (including the problem statement, agent trajectory, and current git diff), generates a score that estimates the probability that the agent has solved the problem. We experiment with using this model to rerank candidate trajectories sampled from a SWE agent policy, and show that such learned verifiers enable effective inference-time scaling for further performance improvement.
5.1.1 Verifier for General-Purpose Prompting
For OpenHands agents (Wang et al., 2024b, c) with general-purpose prompting (§2), we train a verifier (ORM) that takes as input the trajectory $\tau=[o_{1},a_{1},o_{2},a_{2},\dots,o_{n},a_{n}]$, represented as an interleaved sequence of observations and actions, and generates a scalar reward $r\in[0,1]$. Observations $o_{k}$ include the task problem statement, command execution output, error messages, etc; action $a_{k}$ can be bash command or file operations (e.g., edit, view) from the agent.
@@ 5.1.1 Verifier for General-Purpose Prompting
Training and Inference. We fine-tune 32B Qwen2.5-Coder-Instruct to label trajectories as successful or unsuccessful using output tokens <YES> and <NO> respectively.777§B.6 includes the verifier prompt template. For training data, we re-use two sets of trajectories we sampled on SWE-Gym for agent training in §4.2: (1) off-policy trajectories which contain 443 successful trajectories; (2) on-policy trajectories which contain 875 successful trajectories sampled from the fine-tuned Qwen2.5-Coder-Instruct-32B. [^8] We keep only trajectories within 32k-token length for training, which may reduce their number compared to Section 4.2. We combine both on-policy and off-policy trajectories, randomly sample the same amount of unsuccessful trajectories from each subset (1,318 each), and combine them as our dataset for verifier training (total 2,636 trajectories). We fine-tune the model to predict <YES> for successful trajectories and <NO> for unsuccessful ones.
At inference time, conditioned on the prompt and the agent trajectory $\tau$, we use SGLang (Zheng et al., 2024a) to obtain the log probability of the next token being <YES> ($l_{y}$) or <NO> ($l_{n}$). We then calculate the reward as the probability of success by normalizing the log probability: $r=\exp(l_{y})/(\exp(l_{y})+\exp(l_{n}))$.
Metrics. We report two metrics: (1) Pass@$k$, the proportion of tasks with at least one successful solution among $k$ samples, and (2) Best@$k$, the success rate of the highest-reward trajectories selected by the verifier from $k$ samples per task. Pass@$k$ measures solution discovery (upper bound for Best@$k$); Best@$k$ evaluates verifier accuracy. Mean and variance calculation are detailed in §B.1, following Lightman et al. (2023).
@@
Figure 3: Increasing inference-time compute improves performance on SWE-Bench Verified with a learnt verifier. Both the agent and the verifier are a Qwen2.5-Coder-Instruct-32B model fine-tuned on the corresponding dataset (§5.1.1). OpenHands is used as the agent scaffold.
@@
Results. Fig. 3 shows how Pass@$k$ and Best@K scale with the number of sampled agent trajectories using the fine-tuned 32B model as the agent model. Pass@$k$ demonstrates strong improvement, rising from 20.6 to 37.8% resolution rate as $k$ increases from 1 to 8, and up to 42.8@$k$=16. The Best@$k$ metric, which relies on our verifier’s ability to select the best trajectory, demonstrates more modest but steady progress, improving from a resolution rate of 20.6@1 to 29.8@8, and up to 32.0@16. The gap between Pass@$k$ and Best@$k$, due to the imperfect performance of our trained verifier, indicates there is room for improvements in reward modeling for coding agents. Surprisingly, we found that fine-tuning the verifier model using LoRA (Hu et al., 2022) (29.8@8) with Unsloth (Unsloth Team, 2024) performs better than full-parameter fine-tuning (27.2@8), potentially due regularization. Furthermore, as shown in Fig. 1 (bottom), the Best@$k$ curve exhibits strong linearity on a logarithmic scale, indicating a promising scaling behavior.
@@
Training data matters for verifier. We experiment with variations on the choice of training data for our verifier model. Using full-parameter fine-tuning on Qwen-2.5-Coder-Instruct-32B, we use different mixtures of on- and off-policy trajectories, as well as different distributions of successful and unsuccessful trajectories. As shown in Fig. 8, our ablation study demonstrates that the choice of training data can significantly impact verifier performance. Training with a mixture of off-policy and on-policy data yields the best results (our default setting), reaching a resolution rate of 27@8. In contrast, using only on-policy data from the fine-tuned model shows moderate but limited improvement, while training exclusively on off-policy data from Claude and GPT leads to early performance plateaus around 22% resolution rate. Our findings indicate that verifier training benefits most from a diverse dataset combining both off-policy and on-policy examples.
5.1.2 Verifier for Specialized Workflow